# Gaussian-Based Periodic Grand Canonical Density Functional Theory with Implicit Solvation for Computational Electrochemistry

### v2

@@ IV.5 Practical study: silver corrosion
We apply our method to study the corrosion of electrodes in aqueous interfacial chemistry. The surface transformations of electrified metal/solution interfaces are often critical to understanding electrochemical reaction pathways.Gileadi (2011) Recently, the thermodynamics and kinetics of anodic silver atom corrosion from an Ag(100) slab were extensively studied by Kang and coworkers.Kang et al. (2024) Therein, the computations were performed using VASPsol, equivalent to the LPCM method we have implemented.Mathew et al. (2019) To perform constant potential calculations in VASP, one must update the electron number in between SCF cycles until convergence, which can lead to additional computational and manual work.Islam et al. (2023) Due to the body of computational and experimental work available for this system, we believe that this is a well-established benchmark system to demonstrate the value of our method.
Here, we computed the thermodynamics of the corrosion of a single top-layer silver atom on an Ag(100) lattice at various electric biases as described in the work by Kang and coworkers, since undercoordinated sites are usually most susceptible to corrosion. Using our implementation of GCDFT coupled with LPCM, we were able to obtain the potential energy curves for the corrosion of an Ag(100) surface at five different electric biases (Fig. 5). For these calculations, we used the TZVP-MOLOPT-PBE-GTH basis set with the RPBE functional.Hammer, Hansen, and Nørskov (1999) We used a unit cell containing 25 Ag atoms, consisting of a three-layer slab with an additional single-atom layer on top with $\times6$ $\mathbf{k}$-point sampling. We also included two explicit water molecules providing the first coordination shell to the Ag atom coming off the surface. The GPW plane-wave cutoff employed was 4800 eV with a unit-cell height of 60 Å. Due to the slightly slower convergence of the plane wave cutoff at 1.00 V, we report data at 1.00 V with a GPW plane-wave cutoff of 6000 eV. The LPCM solvation model was used with cavitation turned off due to observed numerical instabilities.Kang et al. (2024) The SHE reference used was calibrated to the experimental PZC of Ag(100) (0.609 V).Trasatti and Lust (1999) We note that the reference PZC used in the original work was 0.603 V.
Our calculated equilibrium electrode potential was 0.70 V with an associated activation barrier of 0.20 eV. Both numbers are in good agreement with the numbers reported using VASPsol (0.68 V and 0.21 eV).Kang et al. (2024) All further computational details are included in Appendix D. The computational tools used in this set of calculations can be applied to a wide range of other electrochemical systems.
Figure 5: Potential energy curves for silver corrosion at different potentials relative to the computational SHE. The dashed lines shown are the results reported by Kang and coworkers.Kang et al. (2024) All calculations were performed with the RPBE functional with a GPW plane-wave cutoff of 4800 eV.
@@ V Conclusions
This work presents the implementation of GCDFT with implicit solvation to facilitate computational electrochemistry calculations. We present a novel method to perform such calculations with GTOs. Our method leverages the compactness of the density matrix in a GTO basis, in contrast to a plane wave basis, and performs direct free energy minimization with respect to the density matrix. The implementation of this method is coupled with an implicit solvation model, which accounts for the presence of solvent and electrolytes, consistent with experimental electrochemistry setups. Our method demonstrates accelerated convergence relative to the current state-of-the-art methods for GCDFT. We use our implementation to confirm the results of a silver corrosion study from Kang and coworkers, Kang et al. (2024) which employs VASPsol, Mathew et al. (2019) one of the more widely used implementations for studying electrochemistry applications.
Our method has the advantage of optimizing electron number within an SCF iteration, whereas constant potential calculations in VASPsol require updating electron number in between geometry optimization cycles. We also believe our implementation in a GTO-based code facilitates the usage of hybrid functionals Lee et al. (2021, 2022); Rettig, Lee, and Head-Gordon (2023) and correlated wavefunction methods Gruber et al. (2018); Carbone et al. (2024); Wang and Berkelbach (2021); Ye and Berkelbach (2024); Zen et al. (2016) for more accurate energies, directly benefiting from the entire body of work in quantum chemistry that accelerates these more sophisticated methods. Overall, we believe the methods in this work will set the stage for enabling more realistic computational simulations of complex electrochemical environments with greater predictive power.
@@ Acknowledgements.
This work was funded by the startup funds given by Harvard University. The work of A. Z. N. was also supported by the NSF GRFP Fellowship. The work of A. R. was also supported by the Harvard Quantum Initiative prize postdoctoral fellowship. This work used computational resources from FASRC cluster supported by the Faculty of Arts and Sciences (FAS) Division of Science Research Computing Group at Harvard, and the Delta system at the National Center for Supercomputing Applications through allocation CHE250143 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. We thank Richard Kang for insightful discussions.
@@
Appendix A Derivation of Kohn-Sham Potential due to Solvation
@@
Here, we describe the potential correction that needs to be added to the Kohn-Sham mean-field potential. We rewrite the general energy expression for the solvation energy within linear dielectric and ionic response. The solvation energy expression can be written as
@@
$\lambda_{\text{lq}}(\textbf{r})=-\frac{(\epsilon_{b}-1)}{8\pi}|\nabla\phi(\textbf{r})|^{2}-\frac{\epsilon_{b}\kappa^{2}}{8\pi}\phi(\textbf{r})$ (33)
@@
A.1 LPCM Solvation Model
@@
Now we derive $V_{\text{cav, LPCM}}$. We denote $\partial\rho_{i}$ as the $i$-th component of the spatial gradient of $\rho$, with successive subscripts representing higher order spatial gradients.Andreussi, Dabo, and Marzari (2012)
@@
A.2 CANDLE Solvation Model
We write the CANDLE cavity functions here again, which are obtained from the “solvent-weighted” electronic density $\overline{\rho}=w_{\text{lq}}*\rho$, where $w_{\text{lq}}$ is a Gaussian which has a solvent-dependent width:Sundararaman and Goddard (2015)
@@ A.2 CANDLE Solvation Model
The detailed expressions for cavitation and dispersion have previously been extensively detailed.Sundararaman, Gunceler, and Arias (2014); Sundararaman and Goddard (2015) The cavitation energy can be generalized to the following form, where $\overline{s}$ is the solvent-expanded shape function $\delta_{\sigma_{\text{vdW}}}*s$ where $\sigma_{\text{vdW}}$ is the solvent vdW diameter.
@@
Appendix B Verification of Absolute Vacuum Potential
In traditional periodic electronic structure codes, the electrostatic potential is chosen such that the average of the electrostatic potential is 0. However, the presence of electrolytes in our solvation model sets a natural reference electrostatic potential in the bulk of the electrolyte. In our implementation, we automatically shift this reference potential, so one does not need to add an EFERMI_SHIFT as in a VASPsol calculation.Mathew et al. (2019) We verify our reference potential by plotting the quantity $E-\mu q_{\text{sol}}$, where $\mu$ is the target chemical potential, and $q_{\text{sol}}$ is the solute charge per unit cell at the potential $\mu$. If our reference potential is correct, we should observe a parabolic curve with a maximum at $q_{\text{sol}}=0$.Santos and Schmickler (2004) Shown in Fig A1 is a sample curve for the Ag(100) surface used in the main text.
@@ Appendix B Verification of Absolute Vacuum Potential
Appendix C Fitting Potentials of Zero Charge (PZC)
The raw data for the fitting of TZCs for the PBC and RPBE functionals is in Table 2. All numbers are in eV.
@@ Appendix C Fitting Potentials of Zero Charge (PZC)
Appendix D Computational Details of the Ag Corrosion Calculations
The geometries for all of the calculations are from prior work by Kang and coworkers.Kang et al. (2024) We first converged the GPW plane wave cutoff for our calculations. Shown in Fig. A2 are the results for 1.00 V, which required the largest cutoff to converge fully. We note that the relative energies change minimally when increasing the GPW plane wave density fitting cutoff from 4800 eV to 6000 eV for 1.00 V. In the main text, we use 4800 eV as the GPW plane wave cutoff for 0.00 V, 0.25 V, 0.50 V, and 0.75 V. To ensure that full plane wave cutoff convergence for 1.00 V, 6000 eV was used.

### v3

@@ IV.5 Practical study: silver corrosion
We apply our method to study the corrosion of electrodes in aqueous interfacial chemistry. The surface transformations of electrified metal/solution interfaces are often critical to understanding electrochemical reaction pathways.75 Recently, the thermodynamics and kinetics of anodic silver atom corrosion from an Ag(100) slab were extensively studied by Kang and coworkers.47 Therein, the computations were performed using VASPsol, equivalent to the LPCM method we have implemented.9 To perform constant potential calculations in VASP, one must update the electron number in between SCF cycles until convergence, which can lead to additional computational and manual work.52 Due to the body of computational and experimental work available for this system, we believe that this is a well-established benchmark system to demonstrate the value of our method.
Here, we computed the thermodynamics of the corrosion of a single top-layer silver atom on an Ag(100) lattice at various electric biases as described in the work by Kang and coworkers, since undercoordinated sites are usually most susceptible to corrosion. Using our implementation of GCDFT coupled with LPCM, we were able to obtain the potential energy curves for the corrosion of an Ag(100) surface at five different electric biases (Fig. 5). For these calculations, we used the TZVP-MOLOPT-PBE-GTH basis set with the RPBE functional.74 We used a unit cell containing 25 Ag atoms, consisting of a three-layer slab with an additional single-atom layer on top with $\times6$ $\mathbf{k}$-point sampling. We also included two explicit water molecules providing the first coordination shell to the Ag atom coming off the surface. The GPW plane-wave cutoff employed was 4800 eV with a unit-cell height of 60 Å. Due to the slightly slower convergence of the plane wave cutoff at 1.00 V, we report data at 1.00 V with a GPW plane-wave cutoff of 6000 eV. The LPCM solvation model was used with cavitation turned off due to observed numerical instabilities.47 The SHE reference used was calibrated to the experimental PZC of Ag(100) (0.609 V).73 We note that the reference PZC used in the original work was 0.603 V.
Our calculated equilibrium electrode potential was 0.70 V with an associated activation barrier of 0.20 eV. Both numbers are in good agreement with the numbers reported using VASPsol (0.68 V and 0.21 eV).47 All further computational details are included in the Supporting Information. The computational tools used in this set of calculations can be applied to a wide range of other electrochemical systems.
Figure 5: Potential energy curves for silver corrosion at different potentials relative to the computational SHE. The dashed lines shown are the results reported by Kang and coworkers.47 All calculations were performed with the RPBE functional with a GPW plane-wave cutoff of 4800 eV except for at 1.00 V, where a GPW plane-wave cutoff of 6000 eV was used.
@@ V Conclusions
This work presents the implementation of GCDFT with implicit solvation to facilitate computational electrochemistry calculations. We present a novel method to perform such calculations with GTOs. Our method leverages the compactness of the density matrix in a GTO basis, in contrast to a plane wave basis, and performs direct free energy minimization with respect to the density matrix. The implementation of this method is coupled with an implicit solvation model, which accounts for the presence of solvent and electrolytes, consistent with experimental electrochemistry setups. Our method demonstrates accelerated convergence relative to the current state-of-the-art methods for GCDFT. We use our implementation to confirm the results of a silver corrosion study from Kang and coworkers, 47 which employs VASPsol, 9 one of the more widely used implementations for studying electrochemistry applications.
Our method has the advantage of optimizing electron number within an SCF iteration, whereas constant potential calculations in VASPsol require updating electron number in between geometry optimization cycles. We also believe our implementation in a GTO-based code facilitates the usage of hybrid functionals 60, 61, 62 and correlated wavefunction methods 76, 77, 78, 79, 80, 81 for more accurate energies, directly benefiting from the entire body of work in quantum chemistry that accelerates these more sophisticated methods. Overall, we believe the methods in this work will set the stage for enabling more realistic computational simulations of complex electrochemical environments with greater predictive power.
@@ Acknowledgements.
This work was funded by the startup funds given by Harvard University. The work of A. Z. N. was also supported by the NSF GRFP Fellowship. The work of A. R. was also supported by the Harvard Quantum Initiative prize postdoctoral fellowship. We thank Diptarka Hait for pointing out a typo in Eq. 2 and Richard Kang for pointing out mislabels in Fig. 5. This work used computational resources from FASRC cluster supported by the Faculty of Arts and Sciences (FAS) Division of Science Research Computing Group at Harvard, and the Delta system at the National Center for Supercomputing Applications through allocation CHE250143 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@
Appendix A Details for the Direct Minimization Method
This section will provide more details of our direct minimization method for grand canonical density functional theory. The grand canonical free energy we seek to minimize is
$\Omega=E_{\text{DFT}}-\mu N-\frac{1}{\beta}S_{\text{el}},$ (25)
$E_{\text{DFT}}$ consists of the one-body Hamiltonian energy (h), Coulomb energy (J), and DFT exchange-correlation energy (xc). We discretize the underlying operators with orthonormalized GTOs as defined in the manuscript:
$\frac{\Omega}{N_{k}}=\frac{1}{N_{k}}\sum_{\mathbf{k}}\left(\text{Tr}\!\big(\mathbf{P}^{\mathbf{k}}(\mathbf{h}^{\mathbf{k}}+\tfrac{1}{2}\mathbf{J}^{\mathbf{k}})\big)+E_{\mathrm{xc}}[\rho(\mathbf{r}),\dots]-\mu\,\text{Tr}(\mathbf{P}^{\mathbf{k}})+\frac{1}{\beta}S_{\mathrm{el}}^{\mathbf{k}}\right).$ (26)
As in Mermin’s thermal Hartree-Fock theory, we want to minimize over all density matrices, Hermitian matrices with eigenvalues between 0 and 1.41 To turn this constrained minimization into an unconstrained minimization, we introduce our variational parameter $\{\mathbf{H}\}$, which we define such that $\mathbf{H^{\textbf{k}}}=\mathbf{F^{\textbf{k}}}$ at convergence. At convergence, we know that the electron population follows a Fermi-Dirac distribution, so we introduce $\text{$\Gamma$}^{\mathbf{k}}=\beta\left(\frac{1}{2}\left({\textbf{H}}^{\mathbf{k}}+{\textbf{H}}^{\mathbf{k}H}\right)-\mu\textbf{I}\right)=\textbf{U}^{\mathbf{k}}\text{$\gamma$}^{\mathbf{k}}\textbf{U}^{\mathbf{k}H}$ where U is the unitary matrix which diagonalizes $\Gamma$, and $\gamma$ are the eigenvalues of $\Gamma$. This representation of $\Gamma$ also enforces the Hermiticity of H. We can thus parameterize our density matrix as follows to satisfy the constraint that it must have eigenvalues between 0 and 1:
${\textbf{P}}^{\mathbf{k}}=\frac{1}{\exp(\text{$\Gamma$}^{\mathbf{k}})+1}=\textbf{U}^{\mathbf{k}}\frac{1}{\exp(\text{$\gamma$}^{\mathbf{k}})+1}\textbf{U}^{\mathbf{k}H}=\textbf{U}^{\mathbf{k}}\text{$\rho$}^{\mathbf{k}}\textbf{U}^{\mathbf{k}H}.$ (27)
Now we derive the gradients for direct minimization. To perform direct minimization, we need $\frac{\partial\Omega}{\partial\textbf{H}}$. We do this by differentiating with respect to both the real and imaginary components of H. To do this, we will differentiate with respect to the real and imaginary components of $\Gamma$. These gradients can subsequently be symmetrized to obtain $\frac{\partial\Omega}{\partial\textbf{H}}$ by following the equations in the main text. We will first differentiate with respect to $E_{\text{DFT}}$. Below, we use the fact that $\frac{\partial E_{\text{DFT}}}{\partial\mathbf{P}^{\textbf{k}}}=\mathbf{F}^{\textbf{k}}$.
$\frac{\partial E_{\text{DFT}}}{\partial\Gamma^{\textbf{k}}_{pq}}$ $=\frac{\partial E_{\text{DFT}}}{\partial P_{nm}^{\textbf{k}}}\frac{\partial P_{nm}^{\textbf{k}}}{\partial\Gamma^{\textbf{k}}_{pq}}=F_{mn}^{\textbf{k}}\frac{\partial(U_{mi}^{\textbf{k}}\rho_{i}^{\textbf{k}}U_{ni}^{\textbf{k}*})}{\partial\Gamma_{pq}^{\textbf{k}}}$ (28)
$=F^{\textbf{k}}_{mn}\left(\frac{U_{mi}^{\textbf{k}}}{\partial\Gamma_{pq}^{\textbf{k}}}\rho_{i}^{\textbf{k}}U_{ni}^{\textbf{k}*}+U_{mi}^{\textbf{k}}\rho_{i}^{\textbf{k}}\frac{\partial U_{ni}^{\textbf{k}*}}{\partial\Gamma_{pq}^{\textbf{k}}}+U_{mi}^{\textbf{k}}\frac{\partial\rho_{i}^{\textbf{k}}}{\partial\Gamma_{pq}^{\textbf{k}}}U_{ni}^{\textbf{k}*}\right)$ (29)
We use the fact that $\frac{\partial\rho}{\partial\gamma}=-\rho(1-\rho)$ and the form of the Hermitian eigenvector and eigenvalue derivatives described in the subsequent section to obtain the expression written in the main text:
$\frac{\partial E_{\text{DFT}}}{\partial\text{Re}(\Gamma^{\textbf{k}}_{pq})}$ $={F}^{\textbf{k}}_{mn}\left(U_{pl}^{\textbf{k}*}U^{\textbf{k}}_{qi}U_{nl}^{\textbf{k}}G_{li}^{\textbf{k}}\rho_{i}^{\textbf{k}}U_{mi}^{\textbf{k}*}+U_{ni}^{\textbf{k}}\rho_{i}^{\textbf{k}}U_{pl}^{\textbf{k}}U_{qi}^{\textbf{k}*}U_{ml}^{\textbf{k}*}G_{li}^{\textbf{k}}-U_{ni}^{\textbf{k}}U_{mi}^{\textbf{k}*}\rho_{i}^{\textbf{k}}(1-\rho_{i}^{\textbf{k}})U_{pi}^{\textbf{k}*}U_{qi}^{\textbf{k}}\right)$ (30)
$\frac{\partial E_{\text{DFT}}}{\partial\text{Im}(\Gamma^{\textbf{k}}_{pq})}$ $=i{F}^{\textbf{k}}_{mn}\left(U_{pl}^{\textbf{k}*}U^{\textbf{k}}_{qi}U_{nl}^{\textbf{k}}G_{li}^{\textbf{k}}\rho_{i}^{\textbf{k}}U_{mi}^{\textbf{k}*}-U_{ni}^{\textbf{k}}\rho_{i}^{\textbf{k}}U_{pl}^{\textbf{k}}U_{qi}^{\textbf{k}*}U_{ml}^{\textbf{k}*}G_{li}^{\textbf{k}}-U_{ni}^{\textbf{k}}U_{mi}^{\textbf{k}*}\rho_{i}^{\textbf{k}}(1-\rho_{i}^{\textbf{k}})U_{pi}^{\textbf{k}*}U_{qi}^{\textbf{k}}\right).$ (31)
The remaining two terms of Eq. 25 do not depend on the orbital rotations U, so if we substitute in $\frac{\partial\rho}{\partial\gamma}$ and add back to $\frac{\partial E_{\text{DFT}}}{\partial\text{$\Gamma$}^{\textbf{k}}}$, we get the equations in the main text:
A.1 Eigenvalue and Eigenvector Derivatives of Hermitian Matrices
We first derive a generalized form for eigenvalue and eigenvector derivatives for eigenvalues $\lambda$ and eigenvectors U of some Hermitian matrix H. We denote a singular eigenvector by $u_{i}$, and the derivative with respect to some parameters as $\dot{u}_{i}$, and we define eigenvalues analogously using $\lambda_{i}$. We first derive the general form of the complex eigenvalue and eigenvector derivative assuming only that the eigenvectors are orthonormal; here we define $\textbf{C}:=\textbf{U}^{H}\dot{\textbf{U}}$:
HU $=\textbf{U}\text{$\Lambda$}$
$\dot{\textbf{H}}\textbf{U}-\textbf{U}\dot{\text{$\Lambda$}}$ $=-\textbf{H}\dot{\textbf{U}}+\dot{\textbf{U}}\text{$\Lambda$}$
$\textbf{U}^{H}\dot{\textbf{H}}\textbf{U}-\dot{\text{$\Lambda$}}$ $=\text{$\Lambda$}\textbf{C}+\textbf{C}\text{$\Lambda$}$
$u_{l}^{H}\dot{\textbf{H}}u_{i}-\delta_{li}\dot{\lambda}_{i}$ $=(\lambda_{i}-\lambda_{l})c_{li}$
From the above derivation, we obtain a few facts:
$\dot{\lambda}_{i}$ $=u_{i}^{H}\dot{\textbf{H}}u_{i}$ (34)
$\dot{\textbf{U}}$ $=\textbf{U}\textbf{C}$ (35)
$c_{li}$ $=\frac{u_{l}^{H}\dot{\textbf{H}}u_{i}}{\lambda_{i}-\lambda_{l}}\quad\text{ if $l\neq i$}$ (36)
$\text{Re}(c_{ii})$ $=0$ (37)
When diagonalizing a Hermitian matrix, the eigenvectors are typically only subject to the orthonormality constraint, so they are specified up to some arbitrary phase $\exp(i\theta)$. This nondeterminacy shows up in the derivatives derived above because $\text{Im}(c_{ii})$ is ill-defined. However, because the energy is gauge-invariant, all instances of $c_{ii}$ will necessarily cancel out in the energy derivative, so in our implementation, we set $c_{ii}=0$.
Appendix B Derivation of Kohn-Sham Potential due to Solvation
@@
Here, we describe the potential correction that needs to be added to the Kohn-Sham mean-field potential. We rewrite the general energy expression for the solvation energy in terms of linear dielectric and ionic response. The solvation energy expression can be written as
@@
$\lambda_{\text{lq}}(\textbf{r})=-\frac{(\epsilon_{b}-1)}{8\pi}|\nabla\phi(\textbf{r})|^{2}-\frac{\epsilon_{b}\kappa^{2}}{8\pi}\phi(\textbf{r})^{2}$ (46)
@@
B.1 LPCM Solvation Model
@@
Now we derive $V_{\text{cav, LPCM}}$. We denote $\partial\rho_{i}$ as the $i$-th component of the spatial gradient of $\rho$, with successive subscripts representing higher order spatial gradients.44
@@
B.2 CANDLE Solvation Model
We write the CANDLE cavity functions here again, which are obtained from the “solvent-weighted” electronic density $\overline{\rho}=w_{\text{lq}}*\rho$, where $w_{\text{lq}}$ is a Gaussian which has a solvent-dependent width:43
@@ B.2 CANDLE Solvation Model
The detailed expressions for cavitation and dispersion have previously been extensively detailed.57, 43 The cavitation energy can be generalized to the following form, where $\overline{s}$ is the solvent-expanded shape function $\delta_{\sigma_{\text{vdW}}}*s$ where $\sigma_{\text{vdW}}$ is the solvent vdW diameter.
@@
Appendix C Verification of Absolute Vacuum Potential
In most periodic electronic structure codes, the electrostatic potential is typically chosen such that the average of the electrostatic potential is 0. However, the presence of electrolytes in our solvation model sets a natural reference electrostatic potential in the bulk of the electrolyte. In our implementation, we automatically shift this reference potential, so one does not need to add an EFERMI_SHIFT as in a VASPsol calculation.9 We verify our reference potential by plotting the quantity $E-\mu q_{\text{sol}}$, where $\mu$ is the target chemical potential, and $q_{\text{sol}}$ is the solute charge per unit cell at the potential $\mu$. If our reference potential is correct, we should observe a parabolic curve with a maximum at $q_{\text{sol}}=0$.70 Shown in Fig A1 is a sample curve for the Ag(100) surface used in the main text.
@@ Appendix C Verification of Absolute Vacuum Potential
Appendix D Fitting Potentials of Zero Charge (PZC)
The raw data for the fitting of TZCs for the PBC and RPBE functionals are in Table 2. All numbers are in eV.
@@ Appendix D Fitting Potentials of Zero Charge (PZC)
Appendix E Computational Details of the Ag Corrosion Calculations
The geometries for all of the calculations are from prior work by Kang and coworkers.47 We first converged the GPW plane wave cutoff for our calculations. Shown in Fig. A2 are the results for 1.00 V, which required the largest cutoff to converge fully. We note that the relative energies change minimally when increasing the GPW plane wave density fitting cutoff from 4800 eV to 6000 eV for 1.00 V. In the main text, we use 4800 eV as the GPW plane wave cutoff for 0.00 V, 0.25 V, 0.50 V, and 0.75 V. To ensure full plane wave cutoff convergence for 1.00 V, 6000 eV was used.